bluesea:

how many ounces of trail mix are in a bag that weighs 0.5675 kilograms?(1 pound= 0.454 kg and 1 1 1 pound= 16 ounces) input numeric values.

Mercury:

old question but will respond so this can be closed \[\frac{ 0.5675~kg }{ }\frac{ 1~lb }{ 0.454~kg }\frac{ 16~oz}{ 1~lb }\]